a swimming pool is empty. at 1:00pm a hose is turned on and begins to deliver 5 gallons per minute into swimming pool. thirty minutes later a second hose is turned on and pours 7 gallons per minute into the pool. at what time will there be 1710 gallons of water in the pool. please explain how to work problem thanks
Answer by ankor@dixie-net.com(15652)

a swimming pool is empty.
at 1:00pm a hose is turned on and begins to deliver 5 gallons per minute into swimming pool.
thirty minutes later a second hose is turned on and pours 7 gallons per minute into the pool.
at what time will there be 1710 gallons of water in the pool.
:
Let t = number minutes that both hose are on (12 gal/min starting 1:30)
Find the number of minutes to fill the pool
:
(5*30)+12t = 1710
150 + 12t = 1710
12t = 1710 - 150
12t = 1560
t = 1560/12
t = 130 minutes from 1:30 pm
therefore
1:30 + 2:10 = 3:40 pm will have 1710 gal in the pool